OpenStudy (anonymous):

Find the slope of line passing through the points (11, -2) and (2, -2).

OpenStudy (trojanpoem):

\[slope = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\] slope = -2 - (-2) / -2 -11) = 0